2 years ago I was 4 times my sons age. 8 years ago I was 10 times his age.
How old is he now?

Found 2 solutions by Boreal, macston:
Answer by Boreal(15235)

my age=x; son's age=y
11 years old ANSWER.
x-2=4*(y-2)
x-8=10*(y-8)
x=4y-8+2=4y-6
x=10y-80+8=10y-72
10y-72=4y-6
6y=66
y=11 son's age
two years earlier, son was 9, you were 36
you are 38 now.
eight years ago, you were 30, son was 3.




Answer by macston(5194)

.
Y=your age; S=son's age
.
Y-2=4(S-2)
Y-2=4S-8
Y=4S-6 Use this to substitute for Y below.
.
Y-8=10(S-8)
Y-8=10S-80
Y=10S-72 Substitute for Y from above.
4S-6=10S-72 Subtract 4S from each side.
-6=6S-72 Add 72 to each side.
66=6S
11=S ANSWER: The son is currently 11 years old.
.
CHECK:
Y-2=4(S-2)
Y-2=4(11-2)
Y-2=4(9)
Y-2=36
Y=38 You are currently 38 years old
.
Y-8=10(S-8)
38-8=10(11-8)
30=10(3)
30=30